Grade 12th passAlgebra

If 0.3 the multiplicative inverse of 3 whole 1/3 ? Why or why not?

Harshit Singh
5 Years agoGrade 12th pass

1 Answer

Harshit Singh
5 Years ago
Dear Student

3 whole 1/3 =10/3
0.3 = 3/10
[Multiplicative inverse⟹product should be 1]
According to the question,
3/10×10/3=1
Therefore, 0.3 is the multiplicative inverse of3 whole 1/3

Thanks